OpenStudy (anonymous):

Coordinate Geometry: Find the coordinates of the orthocenter of (triangle symbol)ABC with vertices A(-3, 3), B(-1, 7), and C(3,3). Please help and show work (:!

hero (hero):

Break out the graphing paper

OpenStudy (anonymous):

Can you help me?

OpenStudy (anonymous):

Alright get the general procedure do as directed keep replying

OpenStudy (anonymous):

Find slope of any one side

OpenStudy (anonymous):

Ok, So how would I do that?

OpenStudy (anonymous):

It's better to start with AB: A(-3, 3), B(-1, 7) Do you know slope formula from 2 points?

OpenStudy (anonymous):

y2-y1 and x2 - x1?

OpenStudy (anonymous):

Yep, show us your work by plug the coordinate in, will you!

OpenStudy (anonymous):

m = (y2-y1)/ ( x2 - x1)

OpenStudy (anonymous):

Ok, I get how to do it now

OpenStudy (anonymous):

then find the equation of AB:

OpenStudy (anonymous):

y - yA = m ( x - XA)

OpenStudy (anonymous):

That's the formula?

OpenStudy (anonymous):

Don'i find AB You know slope of AB find slope of line perpendicular as altitude through C is perpendicular to AB find the equation of altitude through C

OpenStudy (anonymous):

Don't we need the intersection point between these segments?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Have you find out slope of AB yet?

OpenStudy (anonymous):

 you have the altitudes